Can you please help me solve the problem...instructions...solve the system by substitution.
3x - 2y = -19
y = 2x + 12

Answer by checkley75(3666)

y=2x+12 then we substitute (2x+12) for y in the other equation and solve for x
3x-2(2x+12)=-19
3x-4x-24=-19
-x=-19+24
-x=5
x=-5 answer now substitute (-5) for x & solve for y
y=2*-5=12
y=-10+12
y=2 answer
proof
3*-5-2*2=-19
-15-4=-19
-19=-19